OpenStudy (anonymous):

The area of a square is represented by 9x2 -42x + 49. Find the length of each side followed by the perimeter. @-------College Al…

OpenStudy (saifoo.khan):

\[(7-3 x)(7-3 x)\]

OpenStudy (saifoo.khan):

\[(7-3 x)^2\]

OpenStudy (saifoo.khan):

Solving for x, 7 -3x = 0 -3x = -7 x = 7/3 <--- length of each side

OpenStudy (anonymous):

@saifoo.khan what would the perimeter be?

OpenStudy (saifoo.khan):

The formula for perimter of square is 4L. 4(7/3) 28/3

OpenStudy (anonymous):

@saifoo.khan THANK YOU!!!!

OpenStudy (saifoo.khan):

@mashburn, No worries! ^_^